# InfoAffect: A Dataset for Affective Analysis of Infographics

### v1

@@ Abstract
Infographics are widely used to convey complex information, yet their affective dimensions remain underexplored due to the scarcity of data resources. We introduce a 3.5k-sample affect‐annotated InfoAffect dataset, which combines textual content with real-world infographics. We first collect the raw data from six domains and aligned them via preprocessing, the accompanied-text–priority method, and three strategies to guarantee the quality and compliance. After that we construct an affect table and use it to constrain annotation. Five state-of-the-art multimodal large language models (MLLMs) then analyze both modalities, and their outputs are fused with Reciprocal Rank Fusion (RRF) algorithm to yield robust affects and confidences. We conducted a user study with two experiments to validate usability and assess InfoAffect dataset using the Composite Affect Consistency Index (CACI), achieving an overall score of 0.986, which indicates high accuracy.
[…]
1 Introduction
Infographics are graphic visual representations of information, data, or knowledge [24, 34]. Their importance has grown in recent years because they present complex data in ways that are both intuitive and visually engaging [43, 24]. These qualities enhance communication and deepen understanding across multiple domains, such as journalism, education, and policy-making, where clear, compelling visual explanations, can materially improve how information is conveyed and comprehended [11, 33, 2]. For instance, journalists are able to craft infographics that visualize global economic disparities [1]. They aggregate indicators into clear comparative visuals, highlight regional contrasts, and surface broad patterns. Readers can grasp the message quickly without wading through lengthy explanatory text. Similarly, in education, well-designed infographics distill difficult knowledge into digestible visual structures [27]. They break down abstract ideas into sequenced elements and concrete cues that support comprehension and recall.
@@ 1 Introduction
Although considerable progress has been made in the analysis of infographics [19, 22, 31], research on the affective dimensions of infographics, such as people‘s emotions, moods, or feelings, remains scarce and underexplored [23]. One approach to advancing affective analysis in infographics is to develop and standardize comprehensive datasets that capture the affective responses elicited by various infographic designs. Previous work [23] compiled a corpus via manual collection and crowdsourced annotations, laying a foundation for subsequent studies in this area. However, the visualization and artificial intelligence research communities currently lack comprehensive dual-modal infographic datasets that incorporate both visual content and textual information. In the absence of such resources, current MLLMs struggle during training to capture the interplay among visual design, textual description, and affects, thereby constraining progress in the affective understanding of infographics.
@@
  Dataset Size Affect Granularity Affect Lable Text Modality Real-world Image Structured Output
ChartGalaxy [25] 1.76M
InfoDet [18] 101K
InfographicVQA [26] 5.5K
Kaggle Infographics Dataset [46] 7.88K
Smile or Scowl? [23] 0.9k Binary Affects
@@
In order to address the limitation that MLLMs perform poorly in the task of understanding the affects in infographics, we introduce InfoAffect dataset. The design of InfoAffect dataset offers three key advantages: First, we curate real-world infographics from six domains rather than synthetics, so that MLLMs trained or fine-tuned on InfoAffect dataset learn from real-world design conventions, linguistic registers, and noise characteristics, thereby achieving stronger downstream generalization. Second, we construct an affect table to guide and constrain affect annotations, ensuring affects remain within a limited label space. This bounding reduces synonym drift and label proliferation. Third, we combine accompanied texts with real-world infographics to mirror how infographics are encountered in practice; this dual-modal design captures the interplay between visual design and linguistic framing and yields more faithful affect predictions than image-only or text-only approaches. which can be used to advance the study of affective analysis. The InfoAffect dataset is available in a public repository at https://github.com/bulichuchu/InfoAffect-dataset In summary, the main contributions are as follows:
  • We introduce the InfoAffect dataset, including 3.5k samples with affect annotations.
  • We introduce a systematic method for extracting affects from the infographics accompanied by textual descriptions, enabling the identification of specific affective expressions.
@@ 2.1 Theories of Affects
The classic affects discrete theory posits the existence of a finite set of affective categories that are universally applicable. Each category is associated with distinct facial features and physiological characteristics [13, 38]. Subsequently, dimensional accounts reconceptualize affect as coordinates in a small number of continuous axes, and affective categories treated as vague regions  [40, 29]. Later, appraisal theories propose that affects arise from evaluative checks of events relative to an individual’s goals, norms, and resources , yielding patterned yet context sensitive response profiles [42]. Over the past decade, constructionist accounts argue that affect categories are not fixed modules but contextually constructed interpretations of core affect using conceptual knowledge and language, predicting substantial intra-category variability [4].
@@
Most recently, Semantic Space Theory (SST) has offered a data-driven alternative to prespecified affect taxonomies [10, 8, 9]. Instead of assuming a fixed set of discrete categories, SST aggregates rich human language, including free descriptions, comparative judgments, and similarity ratings, to embed both felt and perceived affects in a high-dimensional semantic space. In this representation, affects are organized continuously: nearby points indicate fine-grained experiential similarity, and broader regions arise as coherent neighborhoods without being imposed in advance. The geometry captures smooth transitions along affective gradients, preserves contextual nuance through local topology, and admits overlapping groupings that better reflect human judgments. We ground our approach in SST and use its continuous classification and analysis of affects.
2.2 Affect Recognition Methodology
@@
Early approaches to affect recognition relied heavily on unimodal inputs, such as facial-expression classifiers mapping images to affect labels [12] or audio-based sentiment detection [20], typically evaluated in constrained settings with limited variability. However, these early methods struggled to generalize across real-world, multimodal contexts [37], where affect is conveyed jointly through visuals, language, and sound and where distribution shift, occlusion, and noise are common. Intermediate approaches explored different fusion timing and modalities, text, image and audio, with improved robustness in naturalistic stimuli [36]. More recent work therefore adopted fusion strategies that combine multiple modalities to enhance affect‐recognition performance [37], leveraging complementary cues and reducing reliance on any single modal. The latest wave leverages deep learning architectures such as cross-modal transformers and large multimodal pretrained models: they encode textual and visual inputs, apply cross-attention layers to align semantics between modalities, and output affect labels or continuous affect dimensions [21, 49, 16, 48].
@@
In infographic affect analysis, the modality combination is distinctive: visual design and textual description co-exist within a single artifact and jointly shape the viewer’s affect, yet prior methods seldom target this pairing directly. Therefore, we develop a unified pipeline tailored to this setting that combines insights from earlier unimodal and more recent multimodal advances. Rather than committing to a single modeling recipe, the pipeline synthesizes established practices with contemporary architectures under a common formulation. In spirit, it is comprehensive rather than prescriptive: it draws on what the literature has already shown to be effective, accommodates newer techniques as they mature, and provides a consistent basis for interpreting results in terms of both visual design and textual framing. This integration allows us to treat infographic affect as a genuinely multimodal phenomenon while retaining the flexibility needed for diverse tasks, datasets, and evaluation protocols.
2.3 Existing Infographics Datasets
@@
Advances in infographic research have produced a variety of datasets that differ significantly in scale, modality, and supervision. Datasets like ChartGalaxy (1.76M) and InfoDet (101K) focus on visual structure but lack affective labels and textual modalities, making them limited in their ability to capture affects in infographics [25, 18]. InfographicVQA (5.5K) provides structured question–answer pairs for real-world infographics, but it targets visual question answering rather than affective analysis and lacks both affect annotations and a dedicated text modality [26]. The Kaggle Infographics Dataset (7.88K), primarily synthetic or web-scraped, suffers from inconsistent labeling [46], while “Smile or Scowl?” (0.9K) offers binary affect labels but relies solely on the visual modality [23].
@@
To the best of our knowledge, InfoAffect dataset is the only affect-annotated dataset that combines textual content with real-world infographics; although its overall scale is moderate, it offers broad coverage across affective dimensions. This unique combination enables a deeper exploration of how visual design and textual cues interact to shape affective responses, addressing a critical gap in the current infographic research landscape. A side-by-side summary of these datasets is provided in Table 1.
3 Dataset Construction
Figure 1: Overview of the InfoAffect dataset construction process. For raw data, we algin it through three methods, then extract affects with five state-of-the-art MLLMs, each model returns JSON with affective words, and their confidence. Finally fuse results using RRF algorithm.
@@ 3 Dataset Construction
In this section, we describe the full process of InfoAffect dataset construction, from data preparation and affects extraction, through prompt based affect recognition using MLLMs, to formatting and fusion of the final affective outputs, as shown in Fig. 1.
3.1 Data Preparation
@@
The preparation workflow is illustrated in the top panel of Fig. 1. We gathered a large and heterogeneous corpus of 10,000 infographics and their accompanied text from six distinct fields to ensure coverage, diversity, and ecological validity. This topical diversity supports cross-domain evaluation and reduces overfitting to specific content genres, following best practices in dataset and transparency [14, 17]. All data collection was conducted in strict accordance with applicable authorization agreements, robots directives, and platform terms of use. Specifically, the source fields were as follows:
  • Data journalism: Media organizations and online news platforms that regularly publish data-driven reports, often including infographics to provide deeper insights into societal issues;
  • Governmental portals: Official websites that provide publicly accessible reports, often accompanied by infographics to enhance transparency and inform public policy discussions;
@@
  • Non-governmental organizations: Research-driven institutions that publish reports, policy briefs, and white papers, often featuring summary visuals, infographics to communicate findings to a broader audience;
@@
  • Academic and institutional repositories: Digital archives and platforms that host scientific research and educational content, including infographics aimed at making complex research more accessible to the public;
@@
  • Corporate and institutional reports: Annual reports, sustainability reports, and financial statements from companies and organizations, typically released under open access or permissive terms, which include infographics to support business analysis and transparency;
  • Blogs, documentation pages, and social-media posts: Online platforms where experts and influencers share original infographics, technical diagrams, and visual guides related to design, technology trends, and innovations, often facilitating user engagement and education.
@@
From each record, we retained the infographic and all accompanied texts, such as headline, standfirst, caption, or article snippet. After collecting the raw data, we performed several preprocessing steps, all infographics were converted to a canonical RGB format and constrained not to exceed a specified maximum size. We removed corrupted files and filter images below a minimum resolution threshold to avoid unreadable text. To control redundancy, we used perceptual hashing to detect near duplicates [15]. For each duplicate cluster, we kept the image with the highest total pixel count; if there was still a tie, we kept the earliest by timestamp. When infographics were embedded in PDF or HTML, we extracted the focal figure region to support downstream alignment.
@@
After preprocessing, we implemented an accompanied-text–priority method to ensure each infographic was accompanied by faithful textual summaries, we mapped each infographic to a concise, content-faithful textual description that reflected its communicative intent. ❶ Accompanied-text prioritization. If accompanied text was available, we prioritized it as the primary source. ❷ Prompted Description. Otherwise, if accompanied text was absent, we instead used a well-designed prompt to guide GPT-4o to produce the description. This text is concise, factual, and style-consistent.
@@
Finally, we implemented three strategies to guarantee the quality and compliance of the data and text. Each data instance consists of an infographic-text pair, where the image and accompanying text were treated as a unified entity. This structure provided a consistent multimodal representation, which served as the foundation for the subsequent affect analysis.
  • Consistency Check: We compared key facts in the generated text against the corresponding image and flagged any discrepancies for removal or regeneration;
  • Style and Length Check: We verified adherence to journalistic conventions and length constraints, and excluded templated phrasing and speculative markers;
  • Human Review: We conducted manual assessments on stratified samples across sources and topics to validated factual accuracy and the absence of personally identifiable or sensitive information.
3.2 Affect Table
@@
To ensure that the affective output remains within a limited range, we designed an Affect Table as a structured reference for affective extraction. The Affect Table construction process is showed in Fig. 2. We sampled 1,000 infographics from the previously introduced datasets [25, 18, 26, 46, 23], and prompted GPT-4o [35] to extract affects, which were then organized through a two-level classification. Based on this, we constructed the Affect Table. The following paragraphs detail the two-level classification process.
@@
Table 2: Results of the two-level classification. The left column corresponds to the first-level classification results, the middle column contains the second-level classification results, and the right column lists the affects initially generated by GPT-4o.
First-level Second-level Representative Affects
Positive content content, pleased…
hopeful hopeful, optimistic, anticipating…
surprised surprised, amazed, astonished…
proud proud, accomplished, confident…
awestruck awestruck, impressed…
excited excited, enthusiastic, thrilled…
grateful grateful, thankful, appreciative…
Neutral reflective reflective, thoughtful, contemplative, pensive…
calm calm, tranquil, composed, relaxed, at ease…
neutral neutral, even-tempered, matter-of-fact…
attentive attentive, focused, alert, observant…
indifferent indifferent, unconcerned, detached…
uncertain uncertain, unsure, ambivalent…
curious curious, inquisitive, wondering…
Negative bored bored, uninterested, listless…
frustrated frustrated, irritated, thwarted…
shocked shocked, fearful, scared, alarmed, frightened, horrified, terrified, appalled, aghast, afraid…
annoyed annoyed, aggravated, agitated, grouchy, mad, angry…
sad sad, unhappy, depressed, despairing, hopeless, sorrowful, grief, despondent, melancholy, miserable…
concerned concerned, worried, anxious, uneasy, apprehensive, troubled, nervous, distressed, dread…
overwhelmed overwhelmed, overloaded…
@@
The first-level classification aimed to provide an initial categorization of affects, establishing the coarse semantic structure for subsequent fine-grained clustering. While affective states in natural language are continuous and overlapping, most existing affective computing frameworks [32, 5, 10] converge on three dominant affective polarities—positive, neutral, and negative—as fundamental axes of affective evaluation. Following these established frameworks, the affects can be readily divided into three primary categories. However, such a coarse division was clearly insufficient to capture the diversity of affective. Therefore, we conducted a second-level classification within each primary category to achieve a more fine-grained affective representation.
@@
In this phase, this study drew on the SST of affect [10, 8, 9], which posits that affects are continuously distributed in a high-dimensional space. Based on this theory, we projected each word into a point in a high-dimensional affective space using Word2Vec [30]. To identify affective categories from these points, we employed Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) [28], an unsupervised density-based clustering algorithm that did not require pre-specifying the number of clusters. Given a set of affective points $\mathbf{A}=\{\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{n}\}\subset\mathbb{R}^{m}$, where $m$ was the embedding dimension, the algorithm groups pointed into clusters of varying densities, while labeling ambiguous or isolated words as noise. For each affective point $\mathbf{A}_{i}$, we computed its core distance as
@@
InfoAffect dataset exhibited balanced coverage across the three affective polarities and a diverse set of subcategories derived from the SST of affect. As shown in Fig. 3, the hierarchical organization of categories was visualized via a sunburst chart, revealing both broad coverage and structured interrelationships. The balanced radial distribution indicated that the dataset was not overly biased toward any single polarity, while the layered structure supported the internal consistency of the two-level classification and showed that the Affect Table preserves categorical distinctions alongside semantic continuity. The overall classification results were presented in Table 2 and Fig. 3; only the second-level classification formed the final Affect Table. A detailed account of how the model leveraged this table for supervised affect extraction was provided in Sect. 3.3.2.
3.3 Affect Extraction with MLLMs
@@
To perform supervised affect extraction based on the constructed Affect Table, we employed 5 state-of-the-art MLLMs: GPT-4o [35], Claude Sonnet 4 [3], Doubao-1.5-Pro [6], Gemini 2.5 [44], and Qwen3-VL-Plus [39]. These models serve as independent annotators that analyze each infographic–text pair and extract affects guided by the Affect Table. Their outputs were subsequently fused using the RRF algorithm, as described in the following sections.
3.3.1 JSON Format for Affects

### v2

# InfoAffect: Affective Annotations of Infographics in Information Spread

@@ Abstract
Infographics are widely used in social media to convey complex information, yet how they influence users’ affects remains underexplored due to the scarcity of relevant datasets. To address this gap, we introduce a 3.5k-sample affect‐annotated InfoAffect dataset, which combines textual content with real-world infographics. We first collected the raw data from six fields and aligned it via preprocessing, the accompanied-text-priority method, and three strategies to guarantee quality and compliance. After that, we constructed an Affect Table to constrain annotation. We used five state-of-the-art multimodal large language models (MLLMs) to analyze both modalities, and their outputs were fused with Reciprocal Rank Fusion (RRF) algorithm to yield robust affects and confidences. We conducted a user study with two experiments to validate usability and assess InfoAffect dataset using the Composite Affect Consistency Index (CACI), achieving an overall score of 0.608, which indicates high accuracy.The InfoAffect dataset is available in a public repository at https://github.com/bulichuchu/InfoAffect-dataset.
Figure 1: This provides an overview of the InfoAffect dataset, illustrating its dual-modal structure that integrates visual infographics and their accompanied texts. It summarizes the overall workflow, including data collection, preprocessing, affect extraction via MLLMs, and fusion through Reciprocal Rank Fusion.
Introduction
In contemporary social media contexts, information presentation is shifting from text-only formats toward complex multimodal narratives (Amit-Danhi and Shifman 2018). Infographics, which represent information, data, imagery or knowledge (Lankow, Ritchie, and Crooks 2012; Newsom and Haynes 2008), have emerged as a central medium through which social media users acquire complex information, participate in public discussions, and engage in political mobilization (Chang and Buntain 2023). They effectively promote information sharing and engagement across diverse social media users (Chang and Buntain 2023). It could present complex data in ways that are both intuitive and visually engaging (Smiciklas 2012; Lankow, Ritchie, and Crooks 2012). For instance, a recent study (Lee et al. 2022) found that infographic posts on social media outperformed text-only articles in shareability and comprehension, with 90% of users indicating enhanced understanding through visual presentations.
@@ Introduction
Although considerable progress has been made in the analysis of infographics (Jacob 2020; Kunze et al. 2021), research on the affective dimensions of infographics, such as people’s emotions, moods, or feelings, remains scarce and underexplored (Lan et al. 2021). One approach to advancing affective analysis in infographics is to develop datasets that capture the affects elicited by various infographic designs. Previous work (Lan et al. 2021) compiled a corpus via manual collection and crowdsourced annotations, laying a foundation for subsequent studies in this area. However, the social media and visualization research communities currently lack affects annotated dual-modal infographic datasets that incorporate both visual content and textual information. The absence of such datasets limits our ability to systematically study how visual design and textual description jointly shape users’ affects in social media contexts.
@@
  Dataset Size Affect Granularity Affect Label Text Modality Real-world Image Structured Output
ChartGalaxy (Li et al. 2025) 1.76M
InfoDet (Zhu et al. 2025) 101K
InfographicVQA (Mathew et al. 2022) 5.5K
Kaggle Infographics Dataset (Vinodh 2025) 7.88K
Visually29K (Madan et al. 2021) 29K
Smile or Scowl? (Lan et al. 2021) 0.9K Binary Affects
@@
In response to this gap, we introduce InfoAffect dataset. The design of InfoAffect dataset offers three key advantages: First, we curate real-world infographics from six fields rather than synthetics, so that MLLMs trained or fine-tuned on InfoAffect dataset learn from real-world design conventions, linguistic registers, and noise characteristics. Second, we construct an Affect Table to guide and constrain affect annotations, ensuring affects remain within a limited label space. This bounding reduces synonym drift and label proliferation. Third, we combine accompanied texts with real-world infographics to mirror how infographics are encountered in practice; this dual-modal design captures the interplay between visual design and linguistic framing and yields more faithful affect predictions than image-only or text-only approaches. which can be used to advance the study of affective analysis. In summary, the main contributions are as follows:
  • We introduce the InfoAffect dataset, which comprises 3.5k infographic–text pairs, each annotated with affect labels and corresponding confidence scores.
  • We introduce a systematic method for extracting affects from infographic–text pairs, enabling the identification of specific affective expressions.
@@
The classic affects discrete theory posits the existence of a finite set of affective categories that are universally applicable. Each category is associated with distinct features (Ekman 1992; Plutchik 1980). Subsequently, dimensional accounts reconceptualize affect as coordinates in a small number of continuous axes, and affective categories treated as vague regions (Mehrabian 1996). Later, appraisal theories propose that affects arise from evaluative checks of events relative to an individual’s goals, norms, and resources, yielding patterned yet context sensitive response profiles (Scherer 2005). Over the past decade, constructionist accounts argue that affect categories are not fixed modules but contextually constructed interpretations of core affect using conceptual knowledge and language, predicting substantial intra-category variability (Barrett 2017).
@@
However, recent studies have demonstrated limitations of discrete affect models in complex tasks such as music emotion recognition (Eerola and Vuoskoski 2011). Most recently, Semantic Space Theory (SST) has offered a data-driven alternative to prespecified affect taxonomies (Cowen et al. 2019; Cowen and Keltner 2020, 2019). Instead of assuming a fixed set of discrete categories, SST aggregates rich human language, including free descriptions, comparative judgments, and similarity ratings, to embed both felt and perceived affects in a high-dimensional semantic space. In this representation, affects are organized continuously: nearby points indicate fine-grained experiential similarity, and broader regions arise as coherent neighborhoods without being imposed in advance. The geometry captures smooth transitions along affective gradients, preserves contextual nuance through local topology, and admits overlapping groupings that better reflect human judgments. Given the multimodal nature of infographics and their capacity to evoke complex affects, we ground our approach in SST and use its continuous classification and analysis of affects.
Affect Recognition Methodology
@@
Early approaches to affect recognition relied heavily on unimodal inputs, such as facial-expression classifiers mapping images to affect labels (Dino and Abdulrazzaq 2019) or audio-based sentiment detection (Kaushik, Sangwan, and Hansen 2017), typically evaluated in constrained settings with limited variability. However, these early methods struggled to generalize across real-world, multimodal contexts (Pan et al. 2024), where affect is conveyed jointly through visuals, language, and sound and where distribution shift, occlusion, and noise are common. Intermediate approaches explored different fusion timing and modalities, text, image and audio, with improved robustness in naturalistic stimuli (Pan et al. 2023). More recent work therefore adopted fusion strategies that combine multiple modalities to enhance affect‐recognition performance (Pan et al. 2024), leveraging complementary cues and reducing reliance on any single modal. The latest wave leverages deep learning architectures such as large multimodal pretrained models: they encode textual and visual inputs, apply cross-attention layers to align semantics between modalities, and output affect labels or continuous affect dimensions (Ahnert et al. 2025).
@@
In infographic affect analysis, the modality combination is distinctive: visual design and textual description co-exist within a single artifact and jointly shape the viewer’s affect, yet prior methods seldom target this pairing directly for affect extraction tasks (Khatiwada et al. 2025). Therefore, we develop a unified pipeline tailored to this setting that combines insights from earlier unimodal and more recent multimodal advances. Rather than committing to a single modeling recipe, the pipeline synthesizes established practices with contemporary architectures under a common formulation. In spirit, it is comprehensive rather than prescriptive: it draws on what the literature has already shown to be effective, accommodates newer techniques as they mature, and provides a consistent basis for interpreting results in terms of both visual design and textual framing. This integration allows us to treat infographic affect as a genuinely multimodal phenomenon while retaining the flexibility needed for diverse tasks, datasets, and evaluation protocols.
Existing Infographics Datasets
@@
Advances in infographic research have produced a variety of datasets that differ significantly in scale, modality, and supervision. Datasets like ChartGalaxy (1.76M) and InfoDet (101K) focus on visual structure but lack affective labels and textual modalities, making them limited in their ability to capture affects in infographics (Li et al. 2025; Zhu et al. 2025). InfographicVQA (5.5K) provides structured question–answer pairs for real-world infographics, but it targets visual question answering rather than affective analysis and lacks both affect annotations and a dedicated text modality (Mathew et al. 2022). The Kaggle Infographics Dataset (7.88K), primarily synthetic or web-scraped, suffers from inconsistent labeling (Vinodh 2025), while “Smile or Scowl?” (0.9K) offers binary affect labels but relies solely on the visual modality (Lan et al. 2021).
@@
To the best of our knowledge, InfoAffect dataset is the only affect-annotated dataset that combines textual content with real-world infographics; although its overall scale is moderate, it offers broad coverage across affective dimensions. This unique combination enables a deeper exploration of how visual design and textual cues interact to shape affects, addressing a critical gap in the current infographic research landscape. A side-by-side summary of these datasets is provided in Table 1.
Dataset Construction
@@
In this section, we describe the full process of InfoAffect dataset construction, from data preparation and affects extraction, through prompt based affect extraction using MLLMs, to formatting and fusion of the final affective outputs, as shown in the left panel of Figure 1.
Data Preparation
@@
We gathered a large and heterogeneous corpus of 10,000 infographics and their accompanied text from six distinct fields to ensure coverage, diversity, and ecological validity. This topical diversity supports cross-domain evaluation and reduces overfitting to specific content genres, following best practices in dataset and transparency (Gebru et al. 2021a; Hutchinson et al. 2021). All data collection was conducted in strict accordance with applicable authorization agreements, robots directives, and platform terms of use. Specifically, the source fields were as follows:
  • social media posts: Infographics shared on social media platforms like X, such as an infographic illustrating the thermal wind balance (Brandt 2026).
  • Data journalism: Online news platforms that regularly publish data-driven reports, such as Wall Street Journal (Ulick 2015);
  • Governmental portals: Official websites that provide publicly accessible reports, such as Smart city of HK (Technology and Bureau 2025);
@@
  • Non-governmental organizations: Research-driven institutions that publish reports, policy briefs, and white papers, such as ReliefWeb (Office 2025);
@@
  • Academic and institutional repositories: Digital archives and platforms that host scientific research and educational content, including infographics aimed at making complex research more accessible to the public, as discussed in Sect. Existing Infographics Datasets;
@@
  • Corporate and institutional reports: Annual reports, sustainability reports, and financial statements from companies and organizations, typically released under open access or permissive terms, which include infographics to support business analysis and transparency, such as Adobe Digital Trends Report (Adobe. 2025);
@@
From each record, we retained the infographic and all accompanied texts, such as headline, standfirst, caption, or article snippet. After collecting the raw data, we performed several preprocessing steps, all infographics were converted to a canonical RGB format and constrained not to exceed a specified maximum size. We removed corrupted files and filter images below a minimum resolution threshold to avoid unreadable text. To control redundancy, we used perceptual hashing to detect near duplicates (Hao et al. 2021). For each duplicate cluster, we kept the image with the highest total pixel count; if there was still a tie, we kept the earliest by timestamp. When infographics were embedded in PDF or HTML, we extracted the focal figure region to support downstream alignment.
@@
After preprocessing, an accompanied-text–priority method was applied to the remaining 3,500 infographics. This ensures each infographic was accompanied by faithful textual summaries. We mapped each infographic to a concise, content-faithful textual description that reflected its communicative intent. ❶ Accompanied-text prioritization. If accompanied text was available, we prioritized it as the primary source. ❷ Prompted Description. Otherwise, if accompanied text was absent (approximately 11% of the dataset), we instead used a well-designed prompt to guide GPT-4o to produce the description. This text is concise, factual, and style-consistent. The complete prompt is provided in (Appendix A).
@@
Finally, we implemented three strategies to guarantee the quality and compliance of the data and text. Each data instance consists of an infographic-text pair, where the image and accompanying text were treated as a unified entity. This structure provided a consistent multimodal representation, which served as the foundation for the subsequent affect analysis. We performed a consistency check by comparing key facts in the generated text against the corresponding image and flagging any discrepancies for removal or regeneration. We also conducted a style and length check to verify adherence to journalistic conventions and length constraints, and excluded templated phrasing and speculative markers. Additionally, we carried out human review through manual assessments on stratified samples across sources and topics to validate factual accuracy and the absence of personally identifiable or sensitive information.
Affect Table
@@
To ensure that the affective output remains within a limited range, we designed an Affect Table as a structured reference for affective extraction. The Affect Table construction process is showed in Figure 2. We sampled 1,000 infographics from the previously introduced datasets (Li et al. 2025; Zhu et al. 2025; Mathew et al. 2022; Vinodh 2025; Lan et al. 2021), and prompted GPT-4o (OpenAI 2024) to extract affects, which were then organized through a two-level classification. Based on this, we constructed the Affect Table. The following paragraphs detail the two-level classification process.
@@
The first-level classification aimed to provide an initial categorization of affects, establishing the coarse semantic structure for subsequent fine-grained clustering. While affective states in natural language are continuous and overlapping, most existing affective computing frameworks (Mohammad 2022; Buechel and Hahn 2018; Cowen et al. 2019) converge on three dominant affective polarities—positive, neutral, and negative—as fundamental axes of affective evaluation. Following these established frameworks, the affects can be readily divided into three primary categories. However, such a coarse division was insufficient to capture the affective diversity. Therefore, we conducted a second-level classification within each primary category to achieve a more fine-grained representation.
@@
In this phase, this study drew on the SST of affect (Cowen et al. 2019; Cowen and Keltner 2017, 2021), which posits that affects are continuously distributed in a high-dimensional space. Building on this theory, we projected each word into a point in a high-dimensional affective space using Word2Vec (Mikolov et al. 2013). To identify affective categories from these points, we employed Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) (McInnes, Healy, and Astels 2017), an unsupervised density-based clustering algorithm that did not require pre-specifying the number of clusters. Given a set of affective points $\mathbf{A}=\{\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{n}\}\subset\mathbb{R}^{m}$, where $m$ was the embedding dimension, the algorithm groups pointed into clusters of varying densities, while labeling ambiguous or isolated words as noise. For each affective point $\mathbf{A}_{i}$, we computed its core distance as
@@
InfoAffect dataset exhibited balanced coverage across the three affective polarities and a diverse set of subcategories derived from the SST of affect. As shown in Figure 3, the hierarchical organization of categories was visualized via a sunburst chart, revealing both broad coverage and structured interrelationships. The balanced radial distribution indicated that the dataset was not overly biased toward any single polarity, while the layered structure supported the internal consistency of the two-level classification and showed that the Affect Table preserves categorical distinctions alongside semantic continuity. The overall classification results were presented in (Appendix B) and Figure 3; only the second-level classification formed the final Affect Table. A detailed account of how the model leveraged this table for supervised affect extraction was provided in Sect. Prompting Strategies for Affect Extraction.
Affect Extraction with MLLMs
@@
To perform supervised affect extraction based on the constructed Affect Table, we employed 5 state-of-the-art MLLMs: GPT-4o (OpenAI 2024), Claude Sonnet 4 (Anthropic. 2025), Doubao-1.5-Pro (ByteDance 2025), Gemini 2.5 (Team et al. 2025), and Qwen3-VL-Plus (QwenTeam 2025). These models serve as independent annotators that analyze each infographic–text pair and extract affects guided by the Affect Table. Prior research has validated that multi MLLMs annotations can achieve reliability comparable to ground truth (Khatiwada et al. 2025). Their outputs were subsequently fused using the RRF algorithm, as described in the following sections.
JSON Schema for Affects